News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 24 August 2001 7.00am EDT

During the last week, Chandra completed the observing schedule as planned.

Chandra passed through the first three eclipses of the Fall 2001 season with nominal power and thermal performance.

Of note last week was the start of the Cycle 3 cost review following the submission deadline of 14 August. Observations of Cycle 3 targets are expected to begin in the late October timeframe.

The schedule of observations for the coming week is shown below and includes the remaining eclipses of the Fall 2001 season that will end on 30 August.

------------------------------------------
RXJ1532                 ACIS-S       Aug 26
NGC 5055                ACIS-S
3C280                   ACIS-S       Aug 27
Radiation Belts                      Aug 28
A1758                   ACIS-S
RX J1350.0+6007         ACIS-I       Aug 29
A2218                   ACIS-I       Aug 30
Radiation Belts
M31 Nucleus             ACIS-I       Aug 31
SDSS0150+0041           ACIS-S
IRAS00317-2142          ACIS-S
SGR 0526-66             ACIS-S
SGR 0526-66             ACIS-S       Sep 1 
0534-69.9               ACIS-S
NGC 1549                ACIS-S       Sep 2

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports